OpenStudy (anonymous):

What statement makes the open sentence 3x + 20 = 5x + 6 true

OpenStudy (texaschic101):

3x + 20 = 5x + 6 3x - 5x = 6 - 20 -2x = - 14 x = 7 The statement that would nake the open sentence true would be if x = 7 see.. 3(7) + 20 = 5(7) + 6 21 + 20 = 35 + 6 42 = 42 (correct)

OpenStudy (texaschic101):

oops...typo...The statement that would MAKE the open sentence true would be if x = 7